OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 2^-x and y = 4^x + 3 intersect are the solutions of the equation 2^-x = 4^x + 3. (4 points) Part B: Make tables to find the solution to 2^-x = 4^x + 3. Take the integer values of x only between -3 and 3. (4 points) Part C: How can you solve the equation 2^-x = 4^x + 3 graphically? (2 points)

OpenStudy (triciaal):

hint: if a = b and b = c then a= c if 10= 6 + 4 and 10 = 8 + 2 10 = 10 therefore 6 + 4 = 8 + 2 here y = y solve for x

OpenStudy (triciaal):

the intersection is when the (x, y) for one graph is same for the other we already equated the "y's"

OpenStudy (triciaal):

for part B do you know how to do the table?

OpenStudy (anonymous):

ok, and no i don't /:

OpenStudy (triciaal):